OpenStudy (anonymous):

find cos2x and sin2x given sinx=-5/13 and x is in quadrant III

zepdrix (zepdrix):

Hey :)

OpenStudy (anonymous):

hello

zepdrix (zepdrix):

Sorry gimme couple mins +_+ wrappin up a phone call.

zepdrix (zepdrix):

Ok ok ok sorry bout that. Where we at? We gotta evaluate the cos2x and sin2x?

zepdrix (zepdrix):



zepdrix (zepdrix):

So I'm drawing a triangle in quadrant 3.

zepdrix (zepdrix):

We can label the sides like this, since sine = opposite / hypotenuse. Note: Never put the negative sign on the hypotenuse length.

zepdrix (zepdrix):

We can use the Pythagorean Theorem to find the missing leg. What do you get? :D

OpenStudy (anonymous):

12

zepdrix (zepdrix):

Ok good. Since we're in quadrant 3, will it be 12 or -12?

OpenStudy (anonymous):

-12

zepdrix (zepdrix):

Ok good good.

zepdrix (zepdrix):

So we've successfully setup our triangle. Yay! Now let's get to the meat and potatoes of this problem. We need to take our expression,\[\Large\bf\sf \sin2x\]And apply the `Sine Double Angle Identity` to it.

zepdrix (zepdrix):

Remember that one? :)

OpenStudy (anonymous):

would it be -25/78

zepdrix (zepdrix):

Hmmm no :o

zepdrix (zepdrix):

\[\Large\bf\sf \sin2x\quad=\quad 2 \sin x \cos x\quad=\quad 2\left(-\frac{5}{13}\right)\left(-\frac{12}{13}\right)\]

OpenStudy (anonymous):

120/169?

zepdrix (zepdrix):

Mmm ya that looks right! :)

zepdrix (zepdrix):

Do you remember your Cosine Double Angle Identity? You'll need to use that for the other one.